# Doping of a Borexino-like Liquid Scintillator with Tellurium-Diols

### v2

@@ Abstract
One of the most promising approaches for future neutrinoless double beta decay ($0\nu\beta\beta$) searches is to incorporate a candidate isotope into the liquid scintillator of a next-generation neutrino detector. In this study, a sample of the high-performance 1,2,4-Trimethylbenzene-based liquid scintillator from the Borexino detector was loaded with different concentrations of Te-diols. Therefore, a novel and completely water-free synthesis in a non-acidic organic environment at room temperature was used. Key parameters of the loaded samples were analyzed and compared with those of the pure Borexino liquid. Both the emission spectrum and transmission remained nearly unchanged, even at high doping levels. The reduction in light yield was moderate, with approximately 8,400 photons emitted for a 1 MeV energy deposition by an electron at 1% tellurium loading. The time profile of the light emission induced by alpha particles was also investigated, revealing that the scintillation response becomes significantly faster with increasing tellurium concentration.
@@ 1 Introduction
2 Te-diol Synthesis and Doping 1,2,4-Trimethylbenzene-based Liquid Scintillator
For the work described here, the authors were provided with a generous amount of scintillator from the Borexino detector [8] by the Borexino Collaboration. The sample was carefeully stored under protective nitrogen atmosphere in glass bottles and subsequently loaded with Te-diols in the laboratory without any modification. The scintillator is based on the widely-used solvent 1,2,4-Trimethylbenzene (CAS: 95-63-6) known for its excellent energy transfer efficiency and 1.5 g/l 2,5-Diphenyloxazole (CAS: 92-71-7) as fluor [8].
To realize the loading of the 1,2,4-Trimethylbenzene-based LS the techniques described in [7] were reviewed and modified towards a completely water-free synthesis in a non-acidic organic environment. The process enables a Te-diol synthesis at room temperature thus not requiring heating the reactants. An only partially comparable approach was chosen by Ding et al for [9, 10]. For the synthesis of the Te-diols, reported in the present publication, the chemicals as listed in Table 1 were purchased in the highest purity directly available from Sigma-Aldrich and not further purified.
@@ 2 Te-diol Synthesis and Doping 1,2,4-Trimethylbenzene-based Liquid Scintillator
Orthotelluric acid (TeAs) H6TeO6 7803-68-1 purum, $\geq$99.0% Sigma-Aldrich
@@
1-(Dimethylamino)dodecane (DDA) CH3(CH2)11N(CH3)2 112-18-5 $\geq$97.0% Sigma-Aldrich
@@
Table 1: Shown are the chemicals used for the synthesis of the Te-diol compounds. The table lists the linar furmulae as well as CAS numbers. Note that all Chemicals were purchased as pure as directly available from Sigma-Aldrich and therefore not futher purified before the synthesis.
For the synthesis, the orthotelluric acid was first crushed in a mortar to obtain a powder with a fine grain size. The telluric acid was then given into a round flask and neutralized with 1-(dimethylamino)dodecane at room temperature while stirring constantly. Benzene is then added to the mixture. To form the Te-diol complexes in the liquid 1,2-Butanediol is now slowly added using a burette. During this process, water is produced in the reaction, which is removed by constant bubbling the liquid with dried nitrogen. Typically within minutes the milky and cloudy liquid clears up and gets fully colorless and transparent while stirring continuously. For the synthesis the molar ratios of the compounds (similarly to [7]) are 0.5:1:2 (DDA:TeA:BD). A tiny excess of a few milligrams of TeAs ensures that no BD remains, which would have a negative effect on the transparency of the final scintillator, as it is almost insoluble in PC. To remove potential TeAs crystals from the liquid filtration by means of a Büchner funnel equipped with ash-free filter paper (retention: 2 µm) is performed. However, it should be mentioned that only negligible amounts of telluric acid remain in the filter paper. In the final step, the benzene is evaporated at room temperature with constant stirring, leaving behind a colorless, transparent, and highly viscous liquid.
It should be noted here that benzene was deliberately chosen to enable evaporation at room temperature and normal pressure of the organic carrier medium. This ensures the removal of the by-product of the reaction (water) by means of an azeotropic distillation in a simple manner. It is clear that benzene could be replaced by toluene or cyclohexane if necessary, and that azeotropic distillation could be carried out using a rotary evaporator and low pressure at similarly low temperatures. Despite its toxicity, benzene has proven to be ideal and particularly easy to use for the small quantities of Te-diols required here, while toluene should be preferred for later, larger applications.
@@ 3.2 Spectral Transmission of Light
Within the scope of measurement uncertainties, no relevant deterioration in the absorbance of the samples could be observed. Loading with 0.125 % - 2 % of the Te-complex appears to have no relevant influence on the transmission of the Borexino mixture. For reasons of clarity, only the pure scintillator and the sample loaded with 1 % are shown in Figure 2.
@@ 3.3 Light Yields
The amount of scintillation photons produced for a specific energy deposition by a specific particle species, the LS light yield (LY), is a crucial parameter greatly affecting the energy resolution and vertex reconstruction capabilities of a scintillation medium. In a two component scintillator (solvent + fluor) the LY drastically increases with increasing fluor loadings from approximately 1 g/l - 10 g/l. Here, internal losses in the solvent (mainly due to impurities) compete with the energy transfer from the solvent to the fluor. At the critical concentration (typ. <1<1 g/l) for the solvent-fluor combination, about half the maximum light yield is reached [12, 13]. For scintillation cocktails with effective energy transfers between solvent and fluor the critical concentration is lower, leading to higher light yields with lower amounts of fluor.
@@
To evaluate the light yield of the liquid scintillators, the setup shown in Figure 3 and described in detail in [15] was used. A 10.9 ml sample of LS is filled under protective N2 atmosphere into a highly reflective PTFE cell with 2 mm thin UV-transparent glass windows such that the entire volume is free of gas bubbles. Before and after the filling the sample as saturated with protective inert gas by extensively bubbling it with N2. To protect the sample, the cell is sealed under moderate over-pressure. For the readout of the scintillation light, two conventional 1.13-inch PMTs (ETEL9128B) are coupled to the cell windows. During the measurements the sample is irradiated by mono-energetic 662 keV gammas from a 137Cs source (activity: $\sim$370 kBq). The position of the source is thereby fixed with respect to the LS cell. The setup is triggered on a coincidence with a 1.5 inch $\times$ 1.5 inch LaBr3(Ce) detector provided by OST Photonics with 3.05% resolution at 662 keV. In the offline analysis only events with a coincident signal in the 185 keV backscattering pseudo-peak of the crystal detector and thus a 477 keV energy deposition in the liquid scintillator are selected [15, 16].
The measured values for different loadings are shown in Table 2 and graphically depicted in Figure 4. By studying the reproducibility of the light yield measurements for the samples the combined systematic uncertainties were estimated. Each sample was measured 10 times on different days each time freshly filled into the previously cleaned empty cell. The variance of these measurements is interpreted as the influence of systematic effects like HV stability, minor temperature changes in the laboratory, efficiency of the oxygen removal from the LS and other sample preparation and cleaning related differences. By this study the relative systematic uncertainty on the LY of a given sample can be estimated as well as statistical errors.
Figure 3: Schematic drawing of the LY setup. A small $1\times 1$ inch highly reflective PTFE liquid scintillator cell with thin UV-transparent glass windows is filled with the LS sample. Two PMTs are coupled to the cell. The LS is irradiated by monoenergetic gamma-quanta from a 137Cs source. To fix the scattering angle and by that the energy deposition of the gammas in the LS, the setup is triggered in coincidence with a LaBr3(Ce) detector. For further details on the setup and the measurement technique see also [15].
Te Loading 0 % 0.125 % 0.25 % 0.5 % 1.0 % 2.0 %
@@
Figure 4: The plot shows the relative light yield of the sample vs. its Te-concentration. It should be mentioned, that light yields above 104 Photons for electron-like energy depositions are still realized with loadings well exceeding 0.25 %. For loadings exceeding 1.0 % the chemical quenching by then Te-complex molecules severely reduces the light emission in this simple LS baesed on a single solvent (PC) and a single fluor (PPO).
@@ 3.4.1 Light emission time profile setup
Figure 5: Left: Cuvette filled with liquid scintillator before being flushed with nitrogen. At the top center of the cuvette it is possible to see the 244Cm source screwed to the cup. Right: Time profile experimental. In is possible to see: the cuvette on the bottom left, with inserted an optical fiber for the PMT characterization, the HL-PMt on the top left very close to the cell and on the right a neutral optical filter, behind which the $L$-PMT is mounted.
The time response of the setup (IRF: Instrumental Response Function) is characterized in situ by the usage of a diode laser (EPL-405) with a typical pulse width of $\sim$ 70 ps. For the PMTs used in the setup the sharp light pulse can be treated as a delta function and thus enabling a direct characterization of the instrumental response of the LS detector (for further details see [21]). The setup reaches an intrinsic time resolution for the TCSPC of $\sim 0.45$ ns [21].
To further enhance the purity of the time profile samples, the experimental setup is placed inside a muon veto system, composed of two 500 mm × 500 mm × 20 mm EJ-200 plastic scintillator detectors.
@@ 3.4.2 Time Profiles for $\alpha$-Interactions with the Liquid Scintillator
Figure 6: Fluorescence time profile of the different mixtures normalised at maximum. On the left, the whole profile is shown, while on the right, a zoom in the interested region has been plotted. Due to the quenching effect on the light yield, the statistics for high concentrated Tellurium mixture is lower but enough to describe the fluorescence profile in the first hundreds of nanoseconds.
For the time profile study presented in this publication a 244Cm alpha source was deployed into the LS vessel. After extensive flushing with nitrogen gas the data taking took place subsequently. The time profiles for different tellurium loadings of the LS are depicted in Figure 6 in logarithmic scale for both, the full DAQ window of 1600 ns (left plot) and a zoom in (right plot) on the first $\sim$ 100 ns of the light emission decay process.
@@
$F_{fit}(t)=IRF(t-t_{0})\otimes N\sum_{i=1}^{4}\frac{A_{i}}{\tau_{i}-\tau_{r}}\left(e^{-\frac{t-t_{0}}{\tau_{i}}}-e^{-\frac{t-t_{0}}{\tau_{r}}}\right)$ (3.1)
The results of all the fits are shown in table 3, in which it is possible to see the reduction of the decay time components for the all the components, while the fractions A are almost constant.
Te Loading 0 % 0.125 % 0.25 % 0.5 % 1.0 % 2.0 %
$A_{1}$ (%) $68.15\pm 0.15$ $9.17\pm 0.16 66$$6.75\pm 0.31$ $67.47\pm 0.19$ $8.10\pm 0.43 66$$6.86\pm 0.43$
$A_{2}$ (%) $19\pm 0.18 14.14$$14\pm 0.14$ $14.68\pm 0.27$ $14.44\pm 0.21$ $14.68\pm 0.30$ $17.30\pm 0.30$
$A_{3}$ (%) $9.60\pm 0.16$ $9.76\pm 0.11$ $10.12\pm 0.23$ $9.76\pm 0.17$ $9.27\pm 0.26$ $8.79\pm 0.26$
$A_{4}$ (%) $8.20\pm 0.29$ $7.20\pm 0.24$ $8.80\pm 0.48$ $8.60\pm 0.33$ $8.00\pm 0.59$ $7.20\pm 0.59$
$\tau_{1}$ (ns) $3.212\pm 0.001$ $2.789\pm 0.004$ $2.549\pm 0.008$ $2.341\pm 0.001$ $2.010\pm 0.001$ $1.501\pm 0.001$
$\tau_{2}$ (ns) $15.05\pm 0.29$ $2.59\pm 0.20 11$$38\pm 0.35 11.$$1.48\pm 0.25$ $9.62\pm 0.25$ $6.59\pm 0.24$
$\tau_{3}$ (ns) $74.10\pm 2.70$ $67.57\pm 1.12$ $65.36\pm 3.07$ $61.30\pm 2.40$ $55.90\pm 3.20$ $46.92\pm 3.20$
$\tau_{4}$ (ns) $478.9\pm 30.3$ $470.0\pm 30.0$ $82.7\pm 40.2 44$$48.9\pm 30.4$ $385.0\pm 51.6$ $399.8\pm 51.6$
@@ 4 Conclusions
Metal-loaded scintillators will play an important role in the search for neutrino-less double beta decay in the future. Experiments such as SNO+ and KamLAND-Zen have already achieved initial success in this area. In this work, a sample of the scintillator from the Borexino experiment was successfully loaded with Te-diols. To realize this, the techniques of the SNO+ experiment were reviewed and modified towards a completely water-free synthesis in a non-acidic organic environment. The process enables a Te-diol synthesis at room temperature thus not requiring heating the reactants. Key parameters of the loaded sample were examined and compared with the initial sample. The emission spectrum and transmission were virtually unaffected by even high doping factors. The decline in light yield is also moderate, meaning that even with 1% tellurium, approximately 8400 photons for 1 MeV energy deposition by an electron are emitted. The time profile of light emission by alpha particles was also studied. It was shown that the speed of scintillation increases significantly when the tellurium loading is increased. This effect was expected, as the addition of large atoms or molecules to scintillators usually leads to chemical quenching effects that accelerate light emission while reducing the absolute light yield. Overall, it was demonstrated that the loading technique used here is suitable even for pseudocumene-based scintillators, which are known to be very sensitive.

### v3

@@ Abstract
One of the most promising approaches for future neutrinoless double beta decay ($0\nu\beta\beta$) searches is to incorporate a candidate isotope into the liquid scintillator of a next-generation neutrino detector. In this study, a sample of the high-performance 1,2,4-Trimethylbenzene-based liquid scintillator from the Borexino detector was loaded with different concentrations of Te-diols. Therefore, a novel and completely water-free synthesis in a non-acidic organic environment at room temperature was used. Key parameters of the loaded samples were analyzed and compared with those of the pure Borexino liquid scintillator. Both the emission spectrum and optical transmission remained nearly unchanged, even at high doping levels. The reduction in light yield was moderate: approximately 8,400 photons were emitted for a 1 MeV energy deposition by an electron at 1% tellurium-loading. The time profile of the light emission induced by alpha particles was also investigated, revealing that the scintillation response becomes significantly faster with increasing tellurium concentration.
@@ 1 Introduction
2 Te-diol Synthesis and Doping of 1,2,4-Trimethylbenzene-based Liquid Scintillator
For the work described here, the authors were provided with a generous amount of scintillator from the Borexino detector [8] by the Borexino Collaboration. The sample was carefully stored under protective nitrogen atmosphere in glass bottles and subsequently loaded with Te-diols in the laboratory without any modification. The scintillator is based on the widely-used solvent 1,2,4-Trimethylbenzene (CAS: 95-63-6) known for its excellent energy transfer efficiency and 1.5 g/l 2,5-Diphenyloxazole (CAS: 92-71-7) as fluor [8].
To realize the loading of the 1,2,4-Trimethylbenzene-based LS, the techniques described in [7] were reviewed and modified towards a completely water-free synthesis in a non-acidic organic environment. The process enables a Te-diol synthesis at room temperature thus not requiring heating the reactants. A partially comparable approach was reported by Ding et al. for [9, 10]. For the synthesis of the Te-diols, reported in the present publication, the chemicals as listed in Table 1 were purchased in the highest purity directly available from Sigma-Aldrich and not further purified.
@@ 2 Te-diol Synthesis and Doping of 1,2,4-Trimethylbenzene-based Liquid Scintillator
Orthotelluric acid (TeA) H6TeO6 7803-68-1 purum, $\geq$99.0% Sigma-Aldrich
@@
N,N-Dimethyldodecylamine (DDA) CH3(CH2)11N(CH3)2 112-18-5 $\geq$97.0% Sigma-Aldrich
@@
Table 1: Shown are the chemicals used for the synthesis of the Te-diol compounds. The table lists the linear formulae as well as CAS numbers. Note that all chemicals were purchased as pure as directly available from Sigma-Aldrich and therefore not further purified before the synthesis.
For the synthesis, the orthotelluric acid was first ground in a mortar to obtain a fine powder. The telluric acid was then added to a round flask and neutralized with N,N-Dimethyldodecylamine at room temperature while stirring constantly. Benzene was then added to the mixture. To form the Te-diol complexes in the liquid 1,2-Butanediol was slowly added using a burette. During this process, water is produced in the reaction, which is removed by continuously bubbling dried nitrogen through the liquid. Typically within hours the milky and cloudy liquid clears up and becomes fully colorless and transparent while stirring continuously. For the synthesis, the molar ratios of the compounds (similarly to [7]) are 0.5:1:2 (DDA:TeA:BD). A tiny excess of a few milligrams of TeA ensures that no BD remains, which would have a negative effect on the transparency of the final scintillator, as it is almost insoluble in PC. To remove potential TeA crystals from the liquid filtration by means of a Büchner funnel equipped with ash-free filter paper (retention: 2 µm) was performed. However, it should be mentioned that only negligible amounts of telluric acid remain in the filter paper. In the final step, the benzene is evaporated at room temperature with constant stirring, leaving behind a colorless, transparent, and highly viscous liquid.
It should be noted that benzene was deliberately chosen to enable evaporation of the organic carrier medium at room temperature and atmospheric pressure. This ensures the removal of the by-product of the reaction (water) by means of an azeotropic distillation in a simple manner. It is clear that benzene could be replaced by toluene if necessary, and that azeotropic distillation could be carried out using a rotary evaporator and low pressure at similarly low temperatures. Despite its toxicity, benzene has proven to be ideal and particularly easy to use for the small quantities of Te-diols required here, while toluene should be preferred for later, larger applications.
@@ 3.2 Spectral Transmission of Light
Within the scope of measurement uncertainties, no relevant deterioration in the absorbance of the samples could be observed. Loading with 0.125% - 2% of the Te-complex shows no significant influence on the transmission of the Borexino mixture. For reasons of clarity, only the pure scintillator and the sample loaded with 1% are shown in Figure 2.
@@ 3.3 Light Yields
The amount of scintillation photons produced for a specific energy deposition by a specific particle species, the LS light yield (LY), is a crucial parameter greatly affecting the energy resolution and vertex reconstruction capabilities of a scintillation medium. In a two component scintillator (solvent + fluor) the LY drastically increases with increasing fluor loadings from approximately 1 g/l - 10 g/l. Here, internal losses in the solvent (mainly due to impurities) compete with the energy transfer from the solvent to the fluor. At the critical concentration (typ. <1 g/l) for the solvent-fluor combination, about half the maximum light yield is reached [12, 13]. For scintillation cocktails with effective energy transfers between solvent and fluor the critical concentration is lower, leading to higher light yields with lower amounts of fluor.
@@
To evaluate the light yield of the liquid scintillators, the setup shown in Figure 3 and described in detail in [15] was used. A 10.9 ml sample of LS is filled under protective N2 atmosphere into a highly reflective PTFE cell with 2 mm thin UV-transparent glass windows such that the entire volume is free of gas bubbles. Before and after the filling the sample was saturated with inert gas by extensively bubbling N2 through it. To protect the sample, the cell is sealed under moderate over-pressure. For the readout of the scintillation light, two conventional 1.13 inch PMTs (ETEL9128B) are coupled to the cell windows. During the measurements the sample is irradiated by mono-energetic 662 keV gammas from a 137Cs source (activity: $\sim$370 kBq). The position of the source is thereby fixed with respect to the LS cell. The setup is triggered by a coincidence with a $\times\,1.5 inch$ LaBr3(Ce) detector provided by OST Photonics with 3.05% resolution at 662 keV. In the offline analysis only events with a coincident signal in the 185 keV backscattering pseudo-peak of the crystal detector and thus a 477 keV energy deposition in the liquid scintillator are selected [15, 16].
Figure 3: Schematic drawing of the LY setup. A small $1\times 1$ inch highly reflective PTFE liquid scintillator cell with thin UV-transparent glass windows is filled with the LS sample. Two PMTs are coupled to the cell. The LS is irradiated by mono-energetic gamma-quanta from a 137Cs source. To fix the scattering angle and by that the energy deposition of the gammas in the LS, the setup is triggered in coincidence with a LaBr3(Ce) detector. For further details on the setup and the measurement technique see also [15].
The measured values for different loadings are shown in Table 2 and graphically depicted in Figure 4. By studying the reproducibility of the light yield measurements for the samples the combined systematic uncertainties were estimated. Each sample was measured 10 times on different days each time freshly filled into the previously cleaned empty cell. The variance of these measurements is interpreted as the influence of systematic effects like HV stability, minor temperature changes in the laboratory, efficiency of the oxygen removal from the LS and other sample preparation and cleaning-related differences. From this study, both the relative systematic uncertainty and the statistical error on the LY of a given sample can be estimated.
Te-Loading 0% 0.125% 0.25% 0.5% 1.0% 2.0%
@@
Figure 4: The plot shows the relative light yield of the sample vs. its Te-concentration. It should be mentioned, that light yields above 104 Photons for electron-like energy depositions are still realized with loadings well exceeding 0.25%. For loadings exceeding 1.0% the chemical quenching by then Te-complex molecules severely reduces the light emission in this simple LS based on a single solvent (PC) and a single fluor (PPO).
@@ 3.4.1 Light emission time profile setup
Figure 5: Left: Cuvette filled with liquid scintillator before being flushed with nitrogen. At the top center of the cuvette it is possible to see the 244Cm source screwed to the cup. Right: Time profile experimental. It is possible to see: the cuvette on the bottom left, with inserted an optical fiber for the PMT characterization, the HL-PMT on the top left very close to the cell and on the right a neutral optical filter, behind which the $L$-PMT is mounted.
The time response of the setup (IRF: Instrumental Response Function) is characterized in situ by the usage of a diode laser (EPL-405) with a typical pulse width of $\sim$70 ps. For the PMTs used in the setup the sharp light pulse can be treated as a delta function and thus enabling a direct characterization of the instrumental response of the LS detector (for further details see [21]). The setup reaches an intrinsic time resolution for the TCSPC of $\sim$0.45 ns [21]. To further enhance the purity of the time profile samples, the experimental setup is placed inside a muon veto system, composed of two 500 mm × 500 mm × 20 mm EJ-200 plastic scintillator detectors.
@@ 3.4.2 Time Profiles for $\alpha$-interactions with the Liquid Scintillator
For the time profile study presented in this publication a 244Cm alpha source was deployed into the LS vessel. After extensive flushing with nitrogen gas the data taking took place subsequently. The time profiles for different tellurium-loadings of the LS are depicted in Figure 6 in logarithmic scale for both, the full DAQ window of 1600 ns (left plot) and a zoom in (right plot) on the first $\sim$100 ns of the light emission decay process.
Figure 6: Fluorescence time profile of the different mixtures normalised at maximum. On the left, the whole profile is shown, while on the right, a zoom on the interested region has been plotted. Due to the quenching effect on the light yield, the statistics for high concentrated Tellurium mixture is lower but enough to describe the fluorescence profile in the first hundreds of nanoseconds.
@@
$F_{\text{fit}}(t)=IRF(t-t_{0})\otimes N\sum_{i=1}^{4}\frac{A_{i}}{\tau_{i}-\tau_{r}}\left(e^{-\frac{t-t_{0}}{\tau_{i}}}-e^{-\frac{t-t_{0}}{\tau_{r}}}\right)$ (3.1)
The results of all the fits are shown in Table 3, in which it is possible to see the reduction of the decay time components for all the components, while the fractions A are almost constant.
Te-Loading 0% 0.125% 0.25% 0.5% 1.0% 2.0%
$A_{1}$ [%] $68.15\pm 0.15$ $9.17\pm 0.16 66$$6.75\pm 0.31$ $67.47\pm 0.19$ $8.10\pm 0.43 66$$6.86\pm 0.43$
$A_{2}$ [%] $19\pm 0.18 14.14$$14\pm 0.14$ $14.68\pm 0.27$ $14.44\pm 0.21$ $14.68\pm 0.30$ $17.30\pm 0.30$
$A_{3}$ [%] $9.60\pm 0.16$ $9.76\pm 0.11$ $10.12\pm 0.23$ $9.76\pm 0.17$ $9.27\pm 0.26$ $8.79\pm 0.26$
$A_{4}$ [%] $8.20\pm 0.29$ $7.20\pm 0.24$ $8.80\pm 0.48$ $8.60\pm 0.33$ $8.00\pm 0.59$ $7.20\pm 0.59$
$\tau_{1}$ [ns] $3.212\pm 0.001$ $2.789\pm 0.004$ $2.549\pm 0.008$ $2.341\pm 0.001$ $2.010\pm 0.001$ $1.501\pm 0.001$
$\tau_{2}$ [ns] $15.05\pm 0.29$ $2.59\pm 0.20 11$$38\pm 0.35 11.$$1.48\pm 0.25$ $9.62\pm 0.25$ $6.59\pm 0.24$
$\tau_{3}$ [ns] $74.10\pm 2.70$ $67.57\pm 1.12$ $65.36\pm 3.07$ $61.30\pm 2.40$ $55.90\pm 3.20$ $46.92\pm 3.20$
$\tau_{4}$ [ns] $478.9\pm 30.3$ $470.0\pm 30.0$ $82.7\pm 40.2 44$$48.9\pm 30.4$ $385.0\pm 51.6$ $399.8\pm 51.6$
@@ 4 Conclusions
Metal-loaded liquid scintillators are expected to play an important role in the search for neutrino-less double beta decay in the future. Experiments such as SNO+ and KamLAND-Zen have already achieved initial success in this area. In this work, a sample of the scintillator from the Borexino experiment was successfully loaded with Te-diols. To realize this, the techniques of the SNO+ experiment were reviewed and modified towards a completely water-free synthesis in a non-acidic organic environment. The process enables a Te-diol synthesis at room temperature thus not requiring heating the reactants. Key parameters of the loaded sample were examined and compared with the initial sample. The emission spectrum and transmission were virtually unaffected by even high doping factors. The decline in light yield is also moderate, meaning that even with 1% tellurium, approximately 8,400 photons for 1 MeV energy deposition by an electron are emitted. The time profile of light emission by alpha particles was also studied. It was shown that the speed of scintillation increases significantly when the tellurium-loading is increased. This effect was expected, as the addition of large atoms or molecules to scintillators usually leads to chemical quenching effects that accelerate light emission while reducing the absolute light yield. Overall, it was demonstrated that the loading technique used here is suitable even for pseudocumene-based scintillators, which are known to be very sensitive.